3x+4y=5,

Y – 2x = 1

Answer by checkley79(3341)

3x+4y=5 y–2x=1
y=2x+1
replace y in the first equation with(2x+1) & solve for x.
3x+4(2x+1)=5
3x+8x+4=5
11x=5-4
11x=1
x=1/11 ans.
y=2*1/11+1
y=2/11+11/11
y=13/11 ans.
Proof:
3(1/11)+4(13/11)=5
3/11+52/11=5
55/11=5
5=5